Use the elimination method to solve the system of equations
a+8b=3
a+2b=-3

Answer by art123(8)

(a+8b)-(a+2b) =3-(-3)
6b=6
b=1
therefore, a+8b=3
a+8*1=3
a+8=3
a=-5